With a height of 64 in, Nelson was the shortest president of a particular club in the past century. The club presidents of the past century have a mean height of 73.2 in and a standard deviation of 2 2 in. a. What is the positive difference between Nelsons height and the mean? b. How many standard deviations is that [the difference found in part (a)]? c. Convert Nelsons height to a z score d. If we consider usual heights to be those that convert to z scores between -2 and 2, is Nelsons height usual or unusual a. The positive difference between Nelsons height and the mean isin b. The difference is ? standard deviations (Type an integer or decimal rounded to two decimal places as needed.) c. The z score is (Type an integer or decimal rounded to two decimal places as needed) d. Is Nelsons height usual or unusual? O Usual ? Unusual Click to select your answerts)
Answer #1

a) Positive difference = 73.2 - 64 = 9.2

b) Number of standard deviations = 9.2/2.2 = 4.18

c) z - score = -4.18

d) Unusual